# Exact computation of Transfer Entropy with Path Weight Sampling

### v1

@@ Appendix D Choice of reference distribution
Although $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories become overall a complicated nonlinear function of the $X_{j}^{(\nu)}$ trajectory due to the conditional noise sampling, the computations in each step are linear and simple. For accessing conditional distributions of trajectories of other variables, the reference dynamics is worked out similarly. Thus for every marginal probability computation for the $\nu$-th trajectory, a unique reference dynamics is used, which is fine-tuned to that trajectory. This method of choosing a reference dynamics is a numerical analogue of constructing an approximation for the solution to the stochastic filtering equation10, 33, albeit one whose error can be exactly corrected through trajectory reweighting. This drastically reduces the computational cost and makes TE-PWS feasible and accurate.
Appendix E Reduced-variance estimator in jump process
Despite the definition of $\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ in Eq. 11 through both escape and jump propensities, only the latter terms contribute to it on average, as the escape propensities cancel, $\langle\lambda_{ij}\rangle=\langle\lambda_{j}\rangle$2. Thus only the jump propensities at the jumps, marginalized over the conditional distribution of hidden variables, are formally needed. For each marginalization, we need weights of trajectories upto the time of each jump. This is indeed available on-the-fly through the RR scheme. These weights thus provide the conditional distributions $P(X_{l,[0,t_{\alpha}]}|X_{i,[0,t_{\alpha}]},X_{j,[0,t_{\alpha}]})$ and $P(X_{i,[0,t_{\alpha}]},X_{l,[0,t_{\alpha}]}|X_{j,[0,t_{\alpha}]})$, where with a slight abuse of notation, the trajectories are now shown upto jump times $t_{\alpha}$ rather than $k\delta t$. We then use these two conditional distributions to compute $\mathcal{Q}_{ij}$ and $\mathcal{Q}_{j}$ in Eqs. 9 and 10 respectively for the $\alpha$-th jump as,
$\mathcal{Q}_{ij}$ $=\int D[X_{l,[0,t_{\alpha}]}]\;Q_{\alpha}\;P(X_{l,[0,t_{\alpha}]}|X_{i,[0,t_{\alpha}]},X_{j,[0,t_{\alpha}]})$ (24)
$\mathcal{Q}_{j}$ $=\int D[X_{i,[0,t_{\alpha}]}]\;D[X_{l,[0,t_{\alpha}]}]\;Q_{\alpha}\;$
.$\qquad\qquad\qquad.\;P(X_{i,[0,t_{\alpha}]},X_{l,[0,t_{\alpha}]}|X_{j,[0,t_{\alpha}]})$ (25)
where $Q_{\alpha}$ is the jump propensity in the full $d$-dimensional space, which is analytically available. Then using only the jump terms in Eqs. 9-11, we obtain an estimate of the transfer entropy, $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$.
Figure 4: Accuracy of transfer entropy estimates $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$ and $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J}+\mathrm{E})}$ in jump processes. (a) Transfer entropy as a function of time from the jump-based estimate $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$ (red circles), reduced-variance estimate $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J}+\mathrm{E})}$ (blue triangles) and exact estimate from PWS (black dashed line). (b) Error at $T=0.2$, defined as the difference of each estimate from the PWS estimate, as a function of increasing Monte-Carlo averaging. Plotted are errors in the jump-based estimate $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$ (orange squares) and the reduced-variance estimate $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J}+\mathrm{E})}$ (green crosses). Initial conditions for the simulations are $n_{X}=50$ and $n_{Y}=500$, where $n_{X}$ and $n_{Y}$ are the number of species of $X$ and $Y$ respectively. $M_{1}$ for (a) is 25600. $M_{2}=1000$ for both subfigures.
Surprisingly, we find that the statistical error in this estimate can be reduced by an order of magnitude by adding the integral of the escape terms in Eqs. 9 and 10, marginalized with a quadrature of $\delta t$. This arises because fluctuations in the jump terms in Eqs. 9 and 10 are pathwise anti-correlated to those in the escape terms, even though the latter cancel on average, as explained below.
@@ Appendix E Reduced-variance estimator in jump process
$\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ $=\Bigg{\langle}Q_{ij}\delta t\ln\frac{Q^{ij}}{Q^{*}}-(Q_{ij}-Q^{*})\delta t\Bigg{\rangle}$ (26)
@@
We see that the second term fluctuates around a mean of 0. By Taylor expanding the first term upto second order for small values of $(Q_{ij}-Q^{*})$, we find
$\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ $=\Bigg{\langle}(Q_{ij}-Q^{*})\delta t$
$+\frac{\delta t(Q_{ij}-Q^{*})^{2}}{2Q^{*}}+\mathcal{O}\big{(}(Q_{ij}-Q^{*})^{3}\big{)}$
$-(Q_{ij}-Q^{*})\delta t\Bigg{\rangle}$ (27)
$=\Bigg{\langle}\frac{\delta t(Q_{ij}-Q^{*})^{2}}{2Q^{*}}+\mathcal{O}\big{(}(Q_{ij}-Q^{*})^{3}\big{)}\Bigg{\rangle}$ (28)
So fluctuations in the second term of Eq. 26 cancel a part of the fluctuations in the first term.
@@
We numerically demonstrate this effect in a chemical reaction network of two species $X$ and $Y$, consisting of reactions $\phi\rightarrow X,X\rightarrow\phi,X\rightarrow X+Y,Y\rightarrow\phi$, with rates 50, 1, 10 and 10 respectively. Plotted in Figs. 4a and b are the two different transfer entropy estimates and their errors as a function of increasing trajectory duration and increasing statistical averaging respectively. As there is no feedback, the transfer entropy is formally equal to the exact mutual information estimate from PWS13, against which we have compared our results. We find that though both estimates yield unbiased results, the reduced-variance estimate $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J}+\mathrm{E})}$ has an order of magnitude smaller error than the jump-based estimate $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$, for the same computational cost. On the other hand, the $\mathcal{O}(\delta t)$ error in $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J}+\mathrm{E})}$ from the quadrature in marginalization is negligible. $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J}+\mathrm{E})}$ is thus a more accurate estimate of the transfer entropy than $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$. We expect this theoretical result to be tested in experiments in the future, such as using data from neural spike trains34.
Appendix F Variants of transfer entropy
@@
A central advance in TE-PWS is the computation of jump propensities by marginalization over hidden variables. Here we show, in the context of discretized Langevin processes, how the same approach can be used to compute other trajectory-based metrics of directional information transmission beyond Schreiber’s transfer entropy.
@@
Directed information. An alternate measure of information transmission from the trajectory $X_{i,[0,N]}$ to $X_{j,[0,N]}$ is given by directed information 17, defined as
$I[X_{i,[0,N]}\to X_{j,[0,N]}]$
$=\sum_{k=0}^{N-1}I\left(X_{j}(k+1);X_{i,[0,k+1]}|X_{j,[0,k]}\right)$
$=\Bigg{\langle}\ln\frac{\prod_{k}P\left(X_{j}(k+1)|X_{i,[0,k+1]},X_{j,[0,k]}\right)}{\prod_{k}P\left(X_{j}(k+1)|X_{j,[0,k]}\right)}\Bigg{\rangle}$ (29)
@@
By comparing the above definition with that of the transfer entropy in Eq. 2, we see that the directed information incorporates the stepwise information transmission into $X_{j}$ coming from the entire trajectory of $X_{i}$ including the current value. We can bring it to a computable form by taking $X_{i}(k+1)$ out of the conditioning in the numerator,
$I[X_{i,[0,N]}\to X_{j,[0,N]}]$
$=\Bigg{\langle}\ln\frac{\prod_{k}P\left(X_{j}(k+1)|X_{i,[0,k+1]},X_{j,[0,k]}\right)}{\prod_{k}P\left(X_{j}(k+1)|X_{j,[0,k]}\right)}\Bigg{\rangle}$
$=\Bigg{\langle}\ln\prod_{k}P\left(X_{i}(k+1),X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)\Bigg{\rangle}$
$-\Bigg{\langle}\ln\prod_{k}P\left(X_{j}(k+1)|X_{j,[0,k]}\right)\Bigg{\rangle}$
$-\Bigg{\langle}\ln\prod_{k}P\left(X_{i}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)\Bigg{\rangle}$ (30)
@@
Here we note that only two marginalization integrals are actually required for computing the three probabilities in Eq. 30. One is over the conditional distribution $P(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]})$ and the other over $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})$, exactly the same as those sampled in TE-PWS for calculating $\mathcal{T}_{X_{i}\to X_{j}}$. These two distributions can give the probabilities in Eq. 30 as
$P\left(X_{i}(k+1),X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)$
$=\int D[X_{l,[0,k]}]\;P\left(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]}\right)$
.$.\;P\left(X_{i}(k+1),X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$ (31)
$P\left(X_{i}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)$
$=\int D[X_{l,[0,k]}]\;P\left(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]}\right)$
.$.\;P\left(X_{i}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$ (32)
$P\left(X_{j}(k+1)|X_{j,[0,k]}\right)$
$=\int\int D[X_{i,[0,k]}]\;D[X_{l,[0,k]}]\;P\left(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}\right)$
.$.\;P\left(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$ (33)
@@
where, aside from the conditional distributions $P(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]})$ and $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})$, all the other probabilities in the integrals are analytically available. Thus, by keeping track of one additional average over the conditional distributions sampled through the RR scheme, TE-PWS can compute directed information with the same complexity as transfer entropy.
@@
Conditional transfer entropy. The conventional transfer entropy can have a positive value even when there is no direct causal link from the input to the output variable, if information is being causally transmitted through intermediate variables. This motivated the definition of a conditional transfer entropy, also known as a causation entropy, that can measure direct causal links35, 24, 36. For any choice of a third variable $X_{m}$, the conditional transfer entropy is defined as
$\mathcal{T}_{X_{i}\to X_{j}|X_{m}}=\sum_{k=0}^{N-1}I\left(X_{j}(k+1);X_{i,[0,k]}|X_{j,[0,k]},X_{m,[0,k]}\right)$
$=\Bigg{\langle}\ln\frac{\prod_{k}P\left(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{m,[0,k]}\right)}{\prod_{k}P\left(X_{j}(k+1)|X_{j,[0,k]},X_{m,[0,k]}\right)}\Bigg{\rangle}$ (34)
@@
This expectation can be computed similar to Eq. 4. The average is computed in a Monte-Carlo fashion over simulated trajectories of all variables. For each set of trajectories, the numerator and denominator are computed by marginalizing over all other hidden variables $X_{l}$ which exclude $X_{m}$ this time. The optimal reference dynamics should now be chosen to incorporate the effects of the $X_{m}$ trajectory through a frozen field of dirft and diffusion, similar to how the $X_{i}$ and $X_{j}$ trajectories influence the reference dynamics as discussed before. Thus, calculation of each conditional transfer entropy with TE-PWS requires two marginalization integrals, similar to the ordinary transfer entropy.
@@
Filtered transfer entropy. Recently, filtered transfer entropy has been proposed as a way to quantify information transfer in the spirit of filtering theory19. The filtered transfer entropy from $X_{i}$ to $X_{j}$ is defined as

### v2

@@ Appendix D Choice of reference distribution
Although $X_{i}^{(\mu)}$ and $X_{l}^{(\mu)}$ trajectories become overall a complicated nonlinear function of the $X_{j}^{(\nu)}$ trajectory due to the conditional noise sampling, the computations in each step are linear and simple. For accessing conditional distributions of trajectories of other variables, the reference dynamics is worked out similarly. Thus for every marginal probability computation for the $\nu$-th trajectory, a unique reference dynamics is used, which is fine-tuned to that trajectory. This method of choosing a reference dynamics is a numerical analogue of constructing an approximation for the solution to the stochastic filtering equation3, 4, albeit one whose error can be exactly corrected through trajectory reweighting. This drastically reduces the computational cost and makes TE-PWS feasible and accurate.
Appendix E Computing transfer entropy for jump processes
Here we describe in full detail the computation of transfer entropy in jump processes. For clarity, we reproduce Eqs. 9-11 from main text here,
@@ Appendix E Computing transfer entropy for jump processes
[…]
$\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ $=\left\langle\pi_{X_{i}\to X_{j}}-\pi_{X_{j}}\right\rangle$ (S14)
where $\pi_{X_{i}\to X_{j}}$ is a functional of the trajectory segments $X_{i,[k,k+1]}$ and $X_{j,[k,k+1]}$, $\pi_{X_{j}}$ is a functional of only the segment $X_{j,[k,k+1]}$, and $\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ is the transfer entropy from $X_{i}$ to $X_{j}$ over the duration from $k\delta t$ to $(k+1)\delta t$. Here $\alpha$ counts the jumps that change the state of $X_{j}$, $N_{j}$ in number, $\mathcal{Q}_{ij},\lambda_{ij}$ and $\mathcal{Q}_{j},\lambda_{j}$ are jump and escape propensities for $X_{j}$ in the marginal spaces of $(X_{i},X_{j})$ and $(X_{j})$ respectively. The marginal jump propensities $\mathcal{Q}_{ij}(\alpha)$ and $\mathcal{Q}_{j}(\alpha)$ in Eqs. S12 and S13 are abbreviations for the marginal jump propensities $\mathcal{Q}_{ij}(\alpha,t_{\alpha})$ and $\mathcal{Q}_{j}(\alpha,t_{\alpha})$ with $t_{\alpha}$ denoting the time of the $\alpha$-th jump along the $X_{j}$ trajectory. The marginal jump propensities for an arbitrary jump labeled as $\beta$ at an arbitrary time $t$ are defined as
$\mathcal{Q}_{ij}(\beta,t)$ $\equiv\int D[X_{l,[0,t]}]\;Q_{\beta}(t)\;P(X_{l,[0,t]}|X_{i,[0,t]},X_{j,[0,t]})$ (S15)
$\mathcal{Q}_{j}(\beta,t)$ $\equiv\int\int D[X_{i,[0,t]}]\;D[X_{l,[0,t]}]\;Q_{\beta}(t)\;P(X_{i,[0,t]},X_{l,[0,t]}|X_{j,[0,t]})$ (S16)
where $\mathcal{Q}_{\beta}(t)$ is the jump propensity in the full $d$-dimensional space, and, with a slight abuse of notation, $X_{i,[0,t]}$ denotes the trajectory of $X_{i}$ from time 0 to $t$. The marginal escape propensities $\lambda_{ij}(t)$ and $\lambda_{j}(t)$ in Eqs. S12 and S13 are defined by summing marginal jump propensities $\mathcal{Q}_{ij}(\beta,t)$ and $\mathcal{Q}_{j}(\beta,t)$ respectively over all possible jumps $\beta$ that change the state of $X_{j}$,
${\lambda}_{ij}(t)$ $\equiv\sum_{\beta}{\mathcal{Q}}_{ij}(\beta,t)$ (S17)
${\lambda}_{j}(t)$ $\equiv\sum_{\beta}{\mathcal{Q}}_{j}(\beta,t)$ (S18)
Though the transfer entropy $\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ is formally defined using both escape and jump propensity terms in Eq. S14, only the latter terms contribute to the transfer entropy on average, as the escape propensity terms cancel on average, $\langle\lambda_{ij}\rangle=\langle\lambda_{j}\rangle$5. This is because for any $\beta$-th jump at any time $t$, $\langle\mathcal{Q}_{ij}(\beta,t)\rangle=\langle\mathcal{Q}_{j}(\beta,t)\rangle$ where the angular brackets denote an average over all trajectories, as can be seen from Eqs. S15 and S16. Thus only the marginal jump propensities at the specific jump times of $X_{j}$, $\mathcal{Q}_{ij}(\alpha)$ and $\mathcal{Q}_{j}(\alpha)$, are formally needed for computing the transfer entropy5. For computing $\mathcal{Q}_{ij}(\alpha)$ and $\mathcal{Q}_{j}(\alpha)$ by marginalization of the jump propensity in the full $d$-dimensional space, $\mathcal{Q}_{\alpha}(t_{\alpha})$, over conditional distributions, as defined in Eqs. S15 and S16, we need the conditional probabilities of the trajectories up to the time of each jump $t_{\alpha}$. This is indeed available on-the-fly from the trajectory weights in the RR scheme, as explained in the End Matter (EM) of the main text. The only difference from the case of the diffusive processes is that in the current case, trajectory weights are computed on-the-fly not only up to times $k\delta t$ and $(k+1)\delta t$, but also up to all intermediate times $t_{\alpha}$ that represent the jump times in $X_{j}$. These weights reflect the conditional distributions $P(X_{l,[0,t_{\alpha}]}|X_{i,[0,t_{\alpha}]},X_{j,[0,t_{\alpha}]})$ and $P(X_{i,[0,t_{\alpha}]},X_{l,[0,t_{\alpha}]}|X_{j,[0,t_{\alpha}]})$. We then compute $\mathcal{Q}_{ij}$ and $\mathcal{Q}_{j}$ in Eqs. S12 and S13 respectively for the $\alpha$-th jump as weighted averages over the trajectories,
$\mathcal{Q}_{ij}(\alpha)$ $=\int D[X_{l,[0,t_{\alpha}]}]\;Q_{\alpha}(t_{\alpha})\;P(X_{l,[0,t_{\alpha}]}|X_{i,[0,t_{\alpha}]},X_{j,[0,t_{\alpha}]})$ (S19)
$\mathcal{Q}_{j}(\alpha)$ $=\int\int D[X_{i,[0,t_{\alpha}]}]\;D[X_{l,[0,t_{\alpha}]}]\;Q_{\alpha}(t_{\alpha})\;P(X_{i,[0,t_{\alpha}]},X_{l,[0,t_{\alpha}]}|X_{j,[0,t_{\alpha}]})$ (S20)
where $Q_{\alpha}(t_{\alpha})$, the jump propensity in the full $d$-dimensional space, is analytically available. Thus, using only the jump propensity terms in Eqs. S12-S14, we obtain an exact estimate of the transfer entropy, which we call $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$. This approach of the computation of the transfer entropy using only the jump propensities has been recently shown to have a faster convergence over an alternate time-discretized approach for neural spike train data6.
Appendix F Reduced variance estimator in jump processes
Surprisingly, we find that the statistical error in the estimate $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$ can be further reduced by an order of magnitude by adding to it the integral of the escape terms in Eqs. S12 and S13, marginalized with a quadrature of $\delta t$. This reduction arises because fluctuations in the jump terms in Eqs. S12 and S13 are pathwise anti-correlated to those in the escape terms, even though the latter cancel on average, as explained below.
The improved estimate is obtained by marginalizing the escape terms in Eq. S12 and S13 over conditional distributions $P\left(X_{l,[0,k]}\big{|}X_{i,[0,k]},X_{j,[0,k]}\right)$ and $P\left(X_{i,[0,k]},X_{l,[0,k]}\big{|}X_{j,[0,k]}\right)$ respectively, rather than $P\left(X_{l,[0,t]}\big{|}X_{i,[0,t]},X_{j,[0,t]}\right)$ and $P\left(X_{i,[0,t]},X_{l,[0,t]}\big{|}X_{j,[0,t]}\right)$ at every instant of time. Such a quadrature for the marginalization is necessary as the integral in the first terms in Eqs. S12 and S13 cannot be evaluated in closed form. The improved estimate, based on both the jump propensities J and the escape propensities E, is thus computed as
[…]
where $\mathcal{T}^{(\mathrm{J})}_{X_{i}\to X_{j}}$ is the estimate from the previous section based on only the jump propensity terms in Eqs. S12 and S13. In Eq. S21, ${\lambda}_{ij}(t)$ and ${\lambda}_{j}(t)$ are computed as ${\lambda}_{ij}(t)\equiv\sum_{\beta}{\mathcal{Q}}_{ij}(\beta,t)$, ${\lambda}_{j}(t)\equiv\sum_{\beta}{\mathcal{Q}}_{j}(\beta,t)$ with $\mathcal{Q}_{ij}(\beta,t)$ and ${\mathcal{Q}}_{j}(\beta,t)$ computed from Eqs. S15 and S16, respectively, but with
${P}(X_{l,[0,t]}|X_{i,[0,t]},X_{j,[0,t]})$ $\simeq P(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]})\;P_{0}(X_{l,[k,t]}|X_{l,[0,k]})$ (S22)
${P}(X_{i,[0,t]},X_{l,[0,t]}|X_{j,[0,t]})$ $\simeq P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})\;P_{0}(X_{i,[k,t]},X_{l,[k,t]}|X_{i,[0,k]},X_{l,[0,k]})$ (S23)
Here $X_{i,[0,k]}$ and $X_{i,[k,t]}$ denote the $X_{i}$ trajectory between times 0 and $k\delta t$ and between times $k\delta t$ and $t$ respectively; $P_{0}$ is, as before, the reference distribution from which we generate the trajectories of the hidden variables. Hence, we see that the estimator $\mathcal{T}^{(\mathrm{J}+\mathrm{E})}_{X_{i}\to X_{j}}$ is computed from the trajectory weights at time $k\delta t$ as available from the RR scheme, rather than the weight at every instant of time $t$. We refer to this as the quadrature approximation for marginalization, which incurs an $\mathcal{O}(\delta t)$ error.
[…]
$\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ $=\Bigg{\langle}Q_{ij}\delta t\ln\frac{Q^{ij}}{Q^{*}}-(Q_{ij}-Q^{*})\delta t\Bigg{\rangle}$ (S24)
@@ Appendix F Reduced variance estimator in jump processes
where the angular brackets now denote an average over $X_{i}$ trajectories. We see that the second term fluctuates around a mean of 0. By Taylor expanding the first term upto second order in $(Q_{ij}-Q^{*})$, we find
$\Delta\mathcal{T}_{[k,k+1]}^{i\to j}$ $=\Bigg{\langle}(Q_{ij}-Q^{*})\delta t+\frac{\delta t(Q_{ij}-Q^{*})^{2}}{2Q^{*}}+\mathcal{O}\big{(}(Q_{ij}-Q^{*})^{3}\big{)}-(Q_{ij}-Q^{*})\delta t\Bigg{\rangle}$ (S25)
$=\Bigg{\langle}\frac{\delta t(Q_{ij}-Q^{*})^{2}}{2Q^{*}}+\mathcal{O}\big{(}(Q_{ij}-Q^{*})^{3}\big{)}\Bigg{\rangle}$ (S26)
So fluctuations in the second term of Eq. S24 cancel a part of the fluctuations in the first term. This results in a smaller overall variance compared to the case where only the first term of Eq. S24 is used. The above proof also holds in the case of multiple kinds of jumps in $X_{j}$, such as increments and decrements in copy numbers, because the escape propensities $\lambda_{ij}$ and $\lambda_{j}$ are additive over the different kinds of jumps and the above proof works separately for each kind of jump.
@@
We numerically demonstrate this effect in a chemical reaction network of two species $X$ and $Y$, consisting of reactions $\phi\rightarrow X,X\rightarrow\phi,X\rightarrow X+Y,Y\rightarrow\phi$, with rate constants $k_{1}=50$, $k_{-1}=1$, $k_{2}=10$ and $k_{-2}=10$ respectively. Plotted in Figs. S1a and b are the two different transfer entropy estimates and their errors as a function of increasing trajectory duration and increasing statistical averaging respectively. As there is no feedback, the transfer entropy is formally equal to the exact mutual information estimate from PWS1, against which we have compared our results. We find that though both estimates yield unbiased results, the reduced-variance estimate $\mathcal{T}_{X\to Y}^{(\mathrm{J}+\mathrm{E})}$ has an order of magnitude smaller error than the jump-based estimate $\mathcal{T}_{X\to Y}^{(\mathrm{J})}$, for the same computational cost. On the other hand, the $\mathcal{O}(\delta t)$ error in $\mathcal{T}_{X\to Y}^{(\mathrm{J}+\mathrm{E})}$ from the quadrature in marginalization is negligible. $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J}+\mathrm{E})}$ is thus a more accurate estimate of the transfer entropy than $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$. We expect this theoretical result to be tested using experimental data in the future, such as using data from neural spike trains6.
Figure S1: Accuracy of transfer entropy estimates $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J})}$ and $\mathcal{T}_{X_{i}\to X_{j}}^{(\mathrm{J}+\mathrm{E})}$ in jump processes. (a) Transfer entropy as a function of time from the jump-based estimate $\mathcal{T}_{X\to Y}^{(\mathrm{J})}$ (red circles), reduced-variance estimate $\mathcal{T}_{X\to Y}^{(\mathrm{J}+\mathrm{E})}$ (blue triangles) and ground truth estimate from PWS (black dashed line) in the chemical reaction network with two species $X$ and $Y$. (b) Error at trajectory duration $T=0.2/k_{-1}$, defined as the difference of each estimate from the PWS estimate, as a function of increasing Monte-Carlo averaging. Plotted are errors in the jump-based estimate $\mathcal{T}_{X\to Y}^{(\mathrm{J})}$ (orange squares) and the reduced-variance estimate $\mathcal{T}_{X\to Y}^{(\mathrm{J}+\mathrm{E})}$ (green crosses). Initial conditions for the simulations are $n_{X}=50$ and $n_{Y}=500$, where $n_{X}$ and $n_{Y}$ are the number of species of $X$ and $Y$ respectively. $M_{1}$ for (a) is 25600. $M_{2}=1000$ for both subfigures.
Appendix G Variants of transfer entropy
@@
A central advance in TE-PWS is the computation of the probability of a dynamical fluctuation by marginalization over trajectories of hidden variables. Here we show, in the context of discretized Langevin processes, how the same approach can be used to compute other trajectory-based metrics of directional information transmission beyond Schreiber’s transfer entropy.
@@
Directed information. An alternate measure of information transmission from the trajectory $X_{i,[0,N]}$ to $X_{j,[0,N]}$ is given by directed information 7, defined as
$I[X_{i,[0,N]}\to X_{j,[0,N]}]=\sum_{k=0}^{N-1}I\left(X_{j}(k+1);X_{i,[0,k+1]}|X_{j,[0,k]}\right)$ $=\Bigg{\langle}\ln\frac{\prod_{k}P\left(X_{j}(k+1)|X_{i,[0,k+1]},X_{j,[0,k]}\right)}{\prod_{k}P\left(X_{j}(k+1)|X_{j,[0,k]}\right)}\Bigg{\rangle}$ (S27)
@@
By comparing the above definition with that of the transfer entropy in Eq. 2 of the main text, we see that the directed information incorporates the stepwise information transmission into $X_{j}$ coming from the entire trajectory of $X_{i}$ including the current value $X_{i}(k+1)$, while Schreiber’s transfer entropy excludes the flow from the current value of $X_{i}$ (compare Eq. S27 against Eqs. 2 and 4 in the main text). We can bring the directed information to a computable form by taking $X_{i}(k+1)$ out of the conditioning in the numerator,
$I[X_{i,[0,N]}\to X_{j,[0,N]}]$ $=\Bigg{\langle}\ln\frac{\prod_{k}P\left(X_{j}(k+1)|X_{i,[0,k+1]},X_{j,[0,k]}\right)}{\prod_{k}P\left(X_{j}(k+1)|X_{j,[0,k]}\right)}\Bigg{\rangle}$
$=\Bigg{\langle}\ln\prod_{k}P\left(X_{i}(k+1),X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)\Bigg{\rangle}-\Bigg{\langle}\ln\prod_{k}P\left(X_{j}(k+1)|X_{j,[0,k]}\right)\Bigg{\rangle}$
$\qquad\qquad-\Bigg{\langle}\ln\prod_{k}P\left(X_{i}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)\Bigg{\rangle}$ (S28)
@@
Here we note that only two marginalization integrals are actually required for computing the three probabilities in Eq. S28. One is over the conditional distribution $P(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]})$ and the other over $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})$, exactly the same as those sampled in TE-PWS for calculating $\mathcal{T}_{X_{i}\to X_{j}}$, where $X_{l}$ denote all variables except $X_{i}$ and $X_{j}$. These two distributions can give the probabilities in Eq. S28 as
$P\left(X_{i}(k+1),X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)$ $=\int D[X_{l,[0,k]}]\;P\left(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]}\right)$
.$\qquad.\;P\left(X_{i}(k+1),X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$ (S29)
$P\left(X_{i}(k+1)|X_{i,[0,k]},X_{j,[0,k]}\right)$ $=\int D[X_{l,[0,k]}]\;P\left(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]}\right)$
.$\qquad.\;P\left(X_{i}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$ (S30)
$P\left(X_{j}(k+1)|X_{j,[0,k]}\right)$ $=\int\int D[X_{i,[0,k]}]\;D[X_{l,[0,k]}]\;P\left(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]}\right)$
.$\qquad.\;P\left(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{l,[0,k]}\right)$ (S31)
@@
where, aside from the two conditional distributions $P(X_{l,[0,k]}|X_{i,[0,k]},X_{j,[0,k]})$ and $P(X_{i,[0,k]},X_{l,[0,k]}|X_{j,[0,k]})$, all the other probabilities in the integrals are analytically available. Thus, by sampling the two conditional distributions through the RR scheme, computing the three averages in Eqs. S29-S31 as averages over those conditional distributions, plugging them in Eq. S28, and evaluating Eq. S28 as a single Monte-Carlo average, TE-PWS can compute directed information with the same computational cost as transfer entropy.
@@
Conditional transfer entropy. The conventional transfer entropy can have a positive value even when there is no direct causal link from the input to the output variable, when information is being causally transmitted through intermediate variables. This motivated the definition of a conditional transfer entropy, also known as causation entropy, that can measure direct causal links8, 9, 10. For any choice of a third variable $X_{m}$, the conditional transfer entropy is defined as
$\mathcal{T}_{X_{i}\to X_{j}|X_{m}}=\sum_{k=0}^{N-1}I\left(X_{j}(k+1);X_{i,[0,k]}|X_{j,[0,k]},X_{m,[0,k]}\right)=\sum_{k}\Bigg{\langle}\ln\frac{P\left(X_{j}(k+1)|X_{i,[0,k]},X_{j,[0,k]},X_{m,[0,k]}\right)}{P\left(X_{j}(k+1)|X_{j,[0,k]},X_{m,[0,k]}\right)}\Bigg{\rangle}$ (S32)
@@
This expectation can be computed similar to the ordinary transfer entropy in Eq. 4 of the main text. The average is computed in a Monte-Carlo fashion over simulated trajectories of all variables. For each set of trajectories, the numerator and denominator are computed by marginalizing over all other hidden variables $X_{l}$ which exclude $X_{m}$ this time. The optimal reference dynamics should now be chosen to incorporate the effects of the $X_{m}$ trajectory through a frozen field of drift and diffusion resulting from the fixed $X_{m}$ trajectory, similar to how the $X_{i}$ and $X_{j}$ trajectories influence the reference dynamics as discussed above, under Choice of reference distribution. Thus, calculation of each conditional transfer entropy with TE-PWS requires two marginalization integrals, similar to the ordinary transfer entropy.
@@
Filtered transfer entropy. Recently, filtered transfer entropy has been proposed as a way to quantify information transfer in the spirit of filtering theory11. The filtered transfer entropy from $X_{i}$ to $X_{j}$ is defined as